OpenStudy (iwanttogotostanford):

Help :

OpenStudy (iwanttogotostanford):

@jim_thompson5910

OpenStudy (mhchen):

u has components of 5, and 6. -2u would multiply them by -2, so the new components of u is <5*-2,6*-2> = <-10,-12> Same idea for v. Then once you have 2 sets, add the x-variable together, and the y-variable together like: u = <a,b> v = <c,d> u + v = <a+c,b+d>

OpenStudy (iwanttogotostanford):

so it would be 0, 18

jimthompson5910 (jim_thompson5910):

v = <-2, -6> 5*v = 5*<-2, -6> 5v = <5*(-2), 5*(-6)> 5v = ???

OpenStudy (iwanttogotostanford):

@jim_thompson5910 18

jimthompson5910 (jim_thompson5910):

you should get a vector

OpenStudy (iwanttogotostanford):

0,18

OpenStudy (iwanttogotostanford):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

for now I'm only focused on `<5*(-2), 5*(-6)>`

OpenStudy (iwanttogotostanford):

10 and -30

jimthompson5910 (jim_thompson5910):

you're close but one value is incorrect

OpenStudy (iwanttogotostanford):

-10

jimthompson5910 (jim_thompson5910):

<5*(-2), 5*(-6)> = <-10, -30>

jimthompson5910 (jim_thompson5910):

-2u + 5v = -2<5,6> + 5<-2,-6> -2u + 5v = <-10,-12> + <-10, -30> add up the corresponding coordinates

OpenStudy (iwanttogotostanford):

yes

jimthompson5910 (jim_thompson5910):

what would <-10,-12> + <-10, -30> simplify to?